# A Persistence-Driven Edit Distance for Graphs with Abstract Weights

### v3

@@ 2.1.3 Functions in Editable Spaces
Consider an editable space $(X,d_{X},\odot_{X},0_{X})$ and a measure space $(M,\mu)$. Then, the space of functions $L_{1}(M,\mu,X):=\{f:M\rightarrow X\mid\int_{M}d_{X}(0_{X},f(p))d\mu(p)<\infty\}/\sim$ with $\sim$ being the equivalence relationship of functions up to zero measure subsets and $\odot_{W}$ being defined pointwise - is editable. In fact, the function $p\mapsto d_{X}(0,f(p))$ is always non negative, so if properties (P3) and (P4) hold for a fixed $p\in M$, then they hold also for integrals. Call $W:=L_{1}(M,X)$, we verify (P3) as follows:
@@ 2.1.4 Finite Products of Editable Spaces
3 Editable Graphs
Now we introduce the objects we will work with. That is, graphs, possibly directed, with a function associating to each edge a weight in a fixed editable space.
@@ 3 Editable Graphs
A directed graph is made by a couple $G=(V_{G},E_{G})$ with $V$ being the set of its vertices and $E_{G}\subseteq V\times V$ the set of its edges. An undirected graph instead is a couple $G=(V_{G},E_{G})$ with the set of edges being a subset of $V\times V/\operatorname*{Sym}(2)$ where the quotient is over the action of $\operatorname*{Sym}(n)$, the symmetric group over $n$ elements.
Note that any directed graph $(V,E)$ can be turned into an undirected graph by considering its the image of the following map: $E\hookrightarrow V\times V\rightarrow V\times V/\operatorname*{Sym}(2)$.
We use the generic word graph when we don’t want to distinguish between directed and undirected. This means implied that the word “graphs” can be replaced with “directed graphs” or “directed graphs” and any edge can be considered either as $e\in V\times V/\operatorname*{Sym}(2)$ or as $e\in V\times V$. We will point out if there is something noteworthy distinguishing the directed and undirected scenarios. The upcoming definitions should clarify how the reader should interpret this simplification process.
@@
An editable (undirected/directed) graph with weights in $X$ is a (undirected/directed) graph $G=(V,E)$ such that: there is a weight function defined on the edges $\varphi_{G}:E\rightarrow X$ with $X$ being an editable space and $0\notin Im(\varphi_{G})$.
Two editable (undirected/directed) graphs $G=(V_{G},E_{G}),G^{\prime}=(V_{G^{\prime}},E_{G^{\prime}})$ are isomorphic if there is a bijection between the vertices $\eta:V_{G}\rightarrow V_{G^{\prime}}$ giving an isomorphism of (undirected/directed) graphs such that $\varphi_{G}(e)=\varphi_{G^{\prime}}(\eta(e))$, with $\eta((v_{1},v_{2})):=(\eta(v_{1}),\eta(v_{2}))$.
Example 1
Positively weighted undirected graphs and positively weighted directed graphs $G=(V,E)$ induce editable graphs with values in $X=\mathbb{R}_{\geq 0}$.
@@
An editable graphs space $(\mathcal{G},X)$ is the set of editable graphs with weight values in $X$ up to editable graphs isomorphisms.
Definition 7
An editable graph is finite if $V_{G}$ and $E_{G}$ are finite sets.
Given a graph $G=(V,E)$ we indicate with $\mathcal{CW}(G)$ the CW-complex built by considering as many 00-cells as vertices in $V$ and a $1$-cell for every edge in $E$ with the gluing function induced by the extremes of the edge. See Figure 2(a) and Figure 2(b). Note that $\mathcal{CW}(G)$ is well defined (up to homeomorphism) also for undirected graphs: exchanging the extremes of the gluing maps gives an homeomorphic $\mathcal{CW}(G)$.
We have a canonical map $V\hookrightarrow\mathcal{CW}(G)$. With an abuse of notation we will often write $v\in\mathcal{CW}(G)$ meaning the image of $v\in V$ via $V\hookrightarrow\mathcal{CW}(G)$. Similarly there is a map $E\hookrightarrow\mathbf{C}(\mathcal{CW}(G))$ (with $\mathbf{C}(U)$ being the set of closed sets of a topological space $U$), induced by the image of the $1$-cells.
We need to establish the following notation.
Definition 8 (Schelp (1972))
A partial semigroup (PSG) is a set $A$ with a partial binary operation: $+:A^{\prime}\rightarrow A$, with $A^{\prime}\subset A\times A$, which is associative, i.e.:
  1. 1.
    $(x,y+z)\in A^{\prime}$ if and only if $(x+y,z)\in A^{\prime}$
  2. 2.
    if $(x,y+z)\in A^{\prime}$ then $(x,y+z)=(x,y+z)$.
We want to work with particular paths defined on the $CW$-complexes $\mathcal{CW}(G)$. To avoid creating ambiguous notation related to the word “paths” and in analogy with the literature on real trees (Evans, 2008), we give the following definition.
Definition 9
Consider a graph $G=(V,E)$. A segment in $G$ is an embedding $s:[0,1]\rightarrow\mathcal{CW}(G)$ such that $s(0),s(1)\in V\hookrightarrow\mathcal{CW}(G)$ and there is a path on $G$, $\gamma_{s}=\{(v_{0},v_{1}),(v_{1},v_{2}),\ldots,(v_{n-1},v_{n})\}\subset E$ such that: $v_{0}=s(0)$, $v_{n}=s(1)$ and $Im(\gamma_{s}\hookrightarrow\mathbf{C}(\mathcal{CW}(G)))=Im(s)$.
For two segments $s,s^{\prime}$ we have $s\sim s^{\prime}$ if $Im(s)=Im(s^{\prime})$. The internal points of a segment are defined as $s^{o}:=Im(s)-\{s(0),s(1)\}$. We collect all equivalence classes of segments in $G$ in the set $\mathcal{S}(G)$.
Remark 1
We have a map $E\hookrightarrow\mathcal{S}(G)$ induced by the attachment maps of the $1$-cells of the $CW$-complex: $e\mapsto s_{e}\in\mathcal{S}(G)$ so that $s_{e}([0,1])=e$. Note that in case $G$ is an undirected graph, $s(t)\sim s(1-t)$.
Proposition 3
$\mathcal{S}(G)$ can be given a PSG structure via paths composition. Consider $s,s^{\prime}\in\mathcal{S}(G)$ such that:
  • $s(1)=s^{\prime}(0)$;
  • $Im(s)\cap Im(s^{\prime})=s(1)$;
then $s+s^{\prime}:=s^{\prime\prime}:[0,1]\rightarrow\mathcal{CW}(G)$ with $s^{\prime\prime}(t)=s(2t)$ for $t\in[0,1/2]$ and $s^{\prime\prime}(t)=s^{\prime}(2t-1)$ for $t\in[1/2,1]$.
(a) A directed graph.
(b) The CW-complex associated to the graph in Figure 2(a).
(c) The sum of segments by paths composition.
(d) Two different segments $s,s^{\prime}$ with their open stars $\mathcal{N}(s^{o})$ and $\mathcal{N}(s^{\prime o})$. We have $\mathcal{N}(s^{\prime o})=s^{\prime o}$.
Figure 2: Plots related to Section 3.
See also Figure 2(c). Note that, as a consequence of 3, the set $\mathcal{S}(G)$ is almost a directed structure on the space $\mathcal{CW}(G)$ (as defined in Fajstrup et al. (2016)), missing just the constant paths.
We state a couple of observations in the following remark.
Remark 2
$\mathcal{S}(G)$ is generated by $s_{e}\in\mathcal{S}(G)$: every $s\in\mathcal{S}(G)$ can be uniquely written as $\sum_{e\in Im(s)}s_{e}$. Thus, the weight function of any editable (directed) graph $\varphi_{G}$ can be extended to a PSG map $\varphi_{G}:\mathcal{S}(G)\rightarrow X$.
Lastly we define the open star of a subset of the $CW$-complex associated to a graph. This definition is standard for simplicial complexes and we extend it roughly by looking at the open star of the graph projected onto the $CW$-complex.
Definition 10
Given $U\subset\mathcal{CW}(G)$, the open star of $U$, $\mathcal{N}(U)\subset\mathcal{CW}(G)$, is the image under $V\hookrightarrow\mathcal{CW}(G)$ of all the vertices $v\in U$, united with $s_{e}^{o}$ for all the edges $e=(v_{1},v_{2})$ such that $v_{1}\in U$ or $v_{2}\in U$.
Particularly important segments are the ones whose open star in $\mathcal{CW}(G)$ is equal to the (open) segment itself. See also Figure 2(d).
Definition 11
We define $\mathcal{O}(G)\subset\mathcal{S}(G)$ as the set of all segments $s\in\mathcal{S}(G)$ such that $\mathcal{N}(s^{o})=s^{o}$.
4 Edits of Graphs
Now we present the main contribution of the paper.
The approach we follow is to define a distance inspired by Edit Distances (Tai, 1979), but with substantial differences in the edit operations. The philosophy of these distances is to allow certain modifications of the base object, called edits, each being associated to a cost, and to define the distance between two objects as the minimal cost that is needed to transform the first object into the second with a finite sequence of edits. In this way, up to properly setting up a set of edits, one can formalize the deformation of a graph modifying the local information induced by the weight function defined on the edges and the graph structure itself. On top of that, edit distances frequently enjoy some decomposition properties which simplify the calculations (Hong et al., 2017), which are notoriously very heavy (Hein et al., 1995). We also point out that, in literature, graphs are often considered with weights being real numbers and avoid modelling the weight space as we do with the definition of editable spaces.
Given an editable (directed) graphs space $(\mathcal{G},X)$, with $(X,\odot,0_{X})$ editable space, we can define our edits.
  • We call shrinking of an edge a change of the weight function. The new weight function must be equal to the previous one on all edges, apart from the “shrunk” one. In other words, for an edge $e$, this means changing the value $\varphi(e)$ with another non zero value in $X$.
  • A deletion is an edit with which an edge $e$ is removed from the graph. If $e=(v_{1},v_{2})$ is the only edge between $v_{1}$ and $v_{2}$, then we also remove either $v_{1}$ or $v_{2}$. The result is well defined up to (directed) graph isomorphism. The inverse of the deletion is the insertion of an edge along with, potentially, the addition of a novel vertex. We can insert an edge between two existing vertices or at a vertex $v$ specifying the name of the new neighbour of $v$, the neighbours of the newly added vertex (that can be either none, or any portion of the neighbours of $v$). Clearly we always need to define the value of the weight function on the new edge.
  • Lastly, consider $s=\sum s_{e}\in\mathcal{O}(G)$. We can remove all edges and vertices in $s^{o}$ and add to $E_{G}$ the edge $(s[0],s[1])$ with, eventually, the direction coherently induced. The weight of the new edge is already encoded in $w:\mathcal{S}(G)\rightarrow X$ and we have $w(s)=\bigodot_{e\in Im(s)}w(e)$. This transformation is called the ghosting of the vertices in $Im(s)$. Its inverse transformation is called the splitting of an edge $e=(v_{1},v_{2})$ and replaces $e$ with some novel vertices $v^{\prime}_{1},\ldots,v^{\prime}_{n}$ and edges $(v_{1},v^{\prime}_{1}),(v^{\prime}_{1},v^{\prime}_{2})\ldots,(v^{\prime}_{n},v_{2})$ which form a segment in the new graph $G^{\prime}$ such that $s\in\mathcal{O}(G^{\prime})$ with $s(0)=v_{1}$, $s(1)=v_{2}$. Note that any $s_{e}\in\mathcal{O}(G)$ induces a trivial ghosting.
(a) Starting and target weighted graphs.
(b) Deletions in red.
(c) Ghostings in yellow.
(d) Shrinkings in green.
(e) Splittings in yellow.
(f) Insertions in red.
Figure 3: (b)$\rightarrow$(e) form an edit path made from the left graph in Figure 3(a) to the right one. In each figure the edges involved in the editing are highlighted with different colors. In the following plot such vertices return black. This edit path can clearly be represented with a mapping - Section 4.2 - which first deletes all the red vertices in Figure 3(b), then ghosts all the yellow segments in Figure 3(c), then shrinks the green edges in Figure 3(d), split edges by means of the yellow vertices in Figure 3(c) and inserts all the red edges in Figure 3(f).
An editable graph $G$ can be edited to obtain another graph, on which one can apply a new edit to obtain a third graph and so on. One can think of this as composing two edits $e_{0}$, $e_{1}$ which are not defined on the same graph, since the second edit is defined on the already edited graph. This is what we mean by composition of edits. Any finite composition of edits is referred to as an edit path. The notations we use are functional notations, even if the edits are not operators, since an edit is not defined on the whole space of graph but on a single graph. For example $e_{1}\circ e_{0}(G)$ means that $G$ is edited with $e_{0}$, and then $e_{0}(G)$ with $e_{1}$. See Figure 3 for an example of an edit path between two graphs whose weights are edges lenghts - Section 2.1.1.
Remark 3
Even if shrinking and deletions are classical edits, with shrinking being usually referred to as relabeling, the ghosting edit is completely unusual and we are not aware of any previous work employing it. Its definition is largely driven by the TDA perspective presented in the Introduction and it is fundamental for the stability results contained in Pegoraro and Secchi (2024) and Pegoraro (2024a) - dealing with the case of $G$ being a tree.
We also give the following definition.
Definition 12
Graphs are equal up to order $2$ vertices if they become isomorphic after applying a finite number of ghostings or splittings. We write $(G,\varphi_{G})\cong_{2}(G^{\prime},\varphi_{G^{\prime}})$.
12 induces an equivalence relationship which identifies the set of graphs inside $(\mathcal{G},X)$ that we want to treat as equal. We call $(\mathcal{G}_{2},X)$ the space of equivalence classes of graphs in $(\mathcal{G},X)$, equal up to order $2$ vertices.
4.1 Costs of Edit Operations
Now we associate to every edit a cost, that is, a length measure in the space $(\mathcal{G},X)$. The costs of the edit operations are defined as follows:
  • the cost of ghosting/splitting operations is always zero. This is coherent with the properties of editable spaces: consider a ghosting merging two edges separated by an order two vertex, with $x$ being the weight of one edge, and $y$ the cost of the other. The total weight of the two edges is the same before and after the ghosting: $\mid d_{X}(x\odot y,0_{X})-d_{X}(x,0_{X})-d_{X}(y,0_{X})\mid=0$.
Definition 13
Given two graphs $G$ and $G^{\prime}$ in $(\mathcal{G},X)$, define:
  • $\Gamma(G,G^{\prime})$ as the set of all finite edit paths between $G$ and $G^{\prime}$;
  • $cost(\gamma)$ as the sum of the costs of the edits for any $\gamma\in\Gamma(G,G^{\prime})$;
  • the graph edit distance as:
    $d_{E}(G,G^{\prime})=\inf_{\gamma\in\Gamma(G,G^{\prime})}cost(\gamma)$
By definition the triangle inequality and symmetry must hold, but, up to now, this edit distance is intractable; one would have to search for all the possible finite edit paths which connect two graphs in order to find the minimal ones. On top of that, having an edit which is completely “for free”, it is not even obvious that $d_{E}(G,G^{\prime})>0$ for some graphs. However, it is clear that $d_{E}$ induces a pseudo-metric on classes of graphs up to order two vertices.
4.2 Mappings
A mapping between the editable graphs $G=(V_{G},E_{G})$ and $G^{\prime}=(V_{G^{\prime}},E_{G^{\prime}})$ is a set $M\subset\mathcal{S}(G)\times\mathcal{S}(G^{\prime})$ satisfying some constraints. Before listing such properties, we introduce the following notation: (1) given two sets $A,B$, $A-B:=\{a\in A\mid a\notin B\}$ (2) for any $M\subset\mathcal{S}(G)\times\mathcal{S}(G^{\prime})$:
$\mathcal{E}_{M}:=\coprod_{e\subset Im(s),s\in\pi_{G}(M)}e$
with $\pi_{G}:\mathcal{S}(G)\times\mathcal{S}(G^{\prime})\rightarrow\mathcal{S}(G)$ being the projection of the cartesian product. We then have the induced maps $\eta_{G}^{M}:\mathcal{E}_{M}\rightarrow E_{G}$. Lastly, $M$ is a mapping if:
  • (G1)
    the maps $\eta_{G}^{M}$ and $\eta_{G^{\prime}}^{M}$ are injective;
  • (G2)
    if $s\in\pi_{G}(M)$ such that $s\in\mathcal{S}(G)-E_{G}$, then $\mathcal{N}(s^{o})-s^{o}\subset G-\pi_{G}(M)$;
  • (G3)
    if we delete all edges in $E_{G}-Im(\eta_{G}^{M})$ and do all the ghostings induced by $s\in\pi_{G}(M)-E_{G}$ (and analogously for $G^{\prime}$), then $M$ induces a graph isomorphism between the resulting subgraphs of $G$ and $G^{\prime}$.
Let’s try to unveil the meaning of (G1)-(G3). The main idea is that any couple in a mapping $M$ identifies a segment of $G$ and a segment of $G^{\prime}$ which are paired and turned one in the other by the mapping.
Condition (G1) is asking that every edge in $G$ and $G^{\prime}$ is at most within one segment which is coupled with a segment of the other graph. We refer to such edges as coupled edges. The edges which are not coupled will be deleted. Condition (G2) ensures that once we remove all edges which are not coupled, and so which are deleted, then we are in the position to make all the ghostings to turn all the segments $s\in\pi_{G}(M)-E_{G}$ (and $s^{\prime}\in\pi_{G^{\prime}}(M)-E_{G^{\prime}}$), in edges of a graph. Lastly, (G3) forces the couples in $M$ to induce a graph isomorphism between the resulting graphs.
Remark 4
Properties (G1)-(G3) are not canonical properties of mappings, and descend from the fact that we have introduced the ghosting operation, which is a novelty of this work. To better compare properties (G1)-(G3) with more classical edit distances we refer to Section 6, where we report (G1)-(G3) for the special case in which the graph is indeed a tree. Mappings in fact are often used to take advantage of the properties of trees to reduce the computational cost of matching two trees.
In what follows we exploit (G1)-(G3) to parametrize a set of edit paths in the editable graphs space starting from $G$ and ending in $G^{\prime}$, which are collected under the name $\gamma_{M}$. We call:
  • $\gamma_{d}^{G}$ a path made by the deletions to be done in $G$, that is, the edges $E_{G}-Im(\eta_{G}^{M})$, executed in any order. So we obtain the graph $G^{M}_{d}=\gamma_{d}^{G}(G)$, which, instead, is well defined and not depending on the order of the deletions.
  • One then proceeds with ghosting all the segments $s\in\pi_{G}(M)-E_{G}$, in any order, getting a path $\gamma^{G}_{g}$ and the graph $G_{M}:=\gamma^{G}_{g}\circ\gamma^{G}_{d}(G)$.
  • By property (G3), $M$ induces an isomorphism between the two graphs $G^{\prime}_{M}$ (defined in the same way as $G_{M}$, but starting from $G^{\prime}$) and $G_{M}$. So one can shrink the edges of $G_{M}$ onto the edges of $G^{\prime}_{M}$ as described by $M$. The composition of all the shrinkings, executed in any order is an edit path $\gamma_{s}^{G}$.
By construction:
$\gamma_{s}^{G}\circ\gamma_{g}^{G}\circ\gamma_{d}^{G}(G)=G^{\prime}_{M},$
and:
$(\gamma_{d}^{G^{\prime}})^{-1}\circ(\gamma_{g}^{G^{\prime}})^{-1}\circ\gamma_{s}^{G}\circ\gamma_{g}^{G}\circ\gamma_{d}^{G}(G)=G^{\prime}.$
where the inverse of an edit path is thought as the composition of the inverses of the single edit operations, taken in the inverse order.
Lastly, we call $\gamma_{M}$ the set of all possible edit paths:
$(\gamma_{d}^{G^{\prime}})^{-1}\circ(\gamma_{g}^{G^{\prime}})^{-1}\circ\gamma_{s}^{G}\circ\gamma_{g}^{G}\circ\gamma_{d}^{G},$
obtained by changing the order in which the edit operations are executed inside $\gamma_{d}$, $\gamma_{g}$ and $\gamma_{s}$. Observe that, even if $\gamma_{M}$ is a set of paths, its cost is well defined:
$cost(M):=cost(\gamma_{M})=cost(\gamma_{d}^{G})+cost(\gamma_{s}^{G})+cost(\gamma_{d}^{G^{\prime}}).$
See Figure 3 for an example of a mapping between weighted graphs.
[…]
Theorem 1 (Main Theorem)
Given two editable graphs $G$ and $G^{\prime}$, for every finite edit path $\gamma$, there is a mapping $M\in\operatorname*{Mapp}(G,G^{\prime})$ such that $cost(M)\leq cost(\gamma)$.
A first corollary immediately follows.
Corollary 2
For any two finite editable graphs $G$ and $G^{\prime}$, the set $\operatorname*{Mapp}(G,G^{\prime})$ is finite. Thus we have the following well defined pseudo-metric between finite editable graphs:
$d_{E}(G,G^{\prime})=\inf\{cost(\gamma)\mid\gamma\in\Gamma(G,G^{\prime})\}=\min\{cost(M)\mid M\in\operatorname*{Mapp}(G,G^{\prime})\}$
which we will refer to as the edit distance between $G$ and $G^{\prime}$.
A second corollary is obtained observing that, if a mapping has cost equal to zero, then it must contain only ghostings.
Corollary 3
Given $G$ and $G^{\prime}$ editable graphs, $d_{E}(G,G^{\prime})=0$ if and only if $G$ and $G^{\prime}$ are equal up to order $2$ vertices. In other words $d_{E}$ is a metric for graphs considered up to order $2$ vertices.
5 A Discussion on the Use of Graphs and Different Editable Spaces
In this section we report a discussion featuring some key examples which regard editable graphs with values in different editable spaces induced from persistent sets (Carlsson and Mémoli, 2013; Curry, 2018; Curry et al., 2022). Persistent sets are objects of great interest in TDA and represent a (non-equivalent) alternative to persistence modules/diagrams when working with information collected from path connected components of filtrations. We stress again that this is to be intended as a roadmap, the outlining of a general idea and not as a formal and coherent description on how one can treat persistent sets via directed editable graphs. In fact, even in the simplest case, i.e. merge trees, there are non-trivial details which need to be taken care of when bridging between persistent sets and editable graphs (see Section 5.5 and Pegoraro (2024a)). Coherently, we explicitly point out which questions and problems should be assessed by future works.
5.1 Finite Posets, DAGs and Transitive Reductions
In this section, following previous works on graphs and posets, we obtain a graph representation of a poset, which we than employ to represent poset-indexed functors.
Consider a finite partially ordered set $(P,\leq)$. It is well known that one can associate to $P$ a unique (up to directed graph isomorphism) directed acyclic graph (DAG) $G_{P}$ so that $G_{P}$ is in bijection with the objects in $P$ and $(a,b)\in E_{G_{P}}$ if and only if $a\leq b$. Moreover, the graph $G_{P}$ can be simplified without losing ordering information.
Definition 14 (adapted from Aho et al. (1972))
A directed graph $G=(V,E)$ is called transitive if for every segment $s\in\mathcal{S}(G)$, $(s(0),s(1))\in E$. For every directed graph $G=(V,E)$ we define its transitive closure as the graph $G^{T}:=(V,E_{G^{T}})$, with $E_{G^{T}}=\bigcap\{E^{\prime}\subset V\times V\mid E\subset E^{\prime},(V,E^{\prime})\text{ transitive graph}\}$.
@@ 5.1 Finite Posets, DAGs and Transitive Reductions
$E_{G^{t}}:=\bigcap\{E^{\prime}\subset V\times V\mid(V,E^{\prime})^{T}=G^{T}\}.$
Given a finite poset $P$, we call $\mathcal{G}(P)$ the transitive reduction of the associated DAG. See Figure 5 for some examples.
5.2 From Persistent Sets to Editable Graphs
Now, first we associate an editable graph structure to a poset $P$. And then we use this to work with functors of the form $S:P\rightarrow\operatorname*{Sets}$, which we call persistent sets, generalising terminology previously introduced by (Carlsson and Mémoli, 2013; Curry, 2018).
If $P$ is finite, we propose two general ways in which we can associate weights to $\mathcal{G}(P)$, relying just on the structure of $P$:
  1. 1.
    suppose now that the poset $P$ is also endowed with a pseudo-metric (see Burago et al. (2022) for a definition) $d_{P}:P\times P\rightarrow\mathbb{R}_{\geq 0}$. Suppose moreover that any time $a<b$, then $d_{P}(a,b)>0$. Then we can associate to $P$ an editable graph $(\mathcal{G}(P),E_{\mathcal{G}(P)},\varphi_{\mathcal{G}(P)})$, with $\varphi_{\mathcal{G}(P)}:E_{\mathcal{G}(P)}\rightarrow\mathbb{R}_{>0}$ defined as $\varphi_{\mathcal{G}(P)}((a,b))=d_{P}(a,b)$. Note that for any poset $P$ with a monotone increasing function $f:P\rightarrow\mathbb{R}$ we can define $d_{P}(a,b):=\mid f(a)-f(b)\mid$ and this is a pseudo metric which satisfies our requirements since, if $a<b$, then $f(a)<f(b)$;
  2. 2.
    instead, suppose there is a map of posets $f:P\rightarrow X$, with $X$ editable space, and the poset structure on $X$ is as defined in 1. Suppose moreover we are in the situation to apply 1. Then, for every edge $(a,b)\in\mathcal{G}(P)$ there is one and only one $r\in X$ with $f(a)+r=f(b)$. Thus we can define $\varphi_{\mathcal{G}(P)}((a,b))=r$.
As further step consider the case of functors of the form $S:P\rightarrow\operatorname*{Sets}$, with $P$ being a finite poset. We might refer to $P$ as the indexing poset. The functor $S$ can be used to push forward onto its image the partial order we have on $P$. We do so associating to $S$ a poset $D_{S}$, called the display poset of $S$, defined as follows.

### v4

@@ 2.1.3 Functions in Editable Spaces
Consider an editable space $(X,d_{X},\odot_{X},0_{X})$ and a measure space $(M,\mu)$. Then, the space of functions:
$L_{1}(M,\mu,X):=\{f:M\rightarrow X\mid\int_{M}d_{X}(0_{X},f(p))d\mu(p)<\infty\}/\sim,$
with $\sim$ being the equivalence relation of functions up to zero measure subsets and $\odot_{W}$ being defined pointwise - is editable. In fact, the function $p\mapsto d_{X}(0,f(p))$ is always non negative, so if properties (P3) and (P4) hold for a fixed $p\in M$, then they hold also for integrals. Call $W:=L_{1}(M,X)$, we verify (P3) as follows:
@@ 2.1.4 Finite Products of Editable Spaces
3 Dendrograms
Now we introduce the objects we will work with: trees, with a function associating to each edge a weight in a fixed editable space.
@@ 3 Dendrograms
A tree structure $T$ is given by a connected rooted acyclic graph $(V_{T},E_{T})$. We indicate the root of the tree with $r_{T}$. The degree of a vertex $v\in V_{T}$ is the number of edges which have that vertex as one of the endpoints. Any vertex with an edge connecting it to the root is its child and the root is its parent: this is the first step of a recursion which defines the parent and children relation for all vertices in $V_{T}.$ The vertices with no children are called leaves and are collected in the set $L_{T}$. The relation $child<parent$ generates a partial order on $V_{T}$ which induces an orientation on the edges. The edges in $E_{T}$ are identified in the form of ordered pairs $(a,b)$ meaning $a<b$ and $a\rightarrow b$. A subtree of a vertex $v$, called $sub_{T}(v)$, is the tree structure whose set of vertices is $\{x\in V_{T}\mid x\leq v\}$.
A key fact is that given a tree structure $T$, identifying an edge $(v,v^{\prime})$ with its lower vertex $v$, gives a bijection between $V_{T}-\{r_{T}\}$ and $E_{T}$, that is $E_{T}\cong V_{T}-\{r_{T}\}$ as sets. Given this bijection, we often use $E_{T}$ to indicate the vertices $v\in V_{T}-\{r_{T}\}$, to simplify the notation.
To avoid overlapping some terminology with already existing notions of trees and for ease of notation, we refer to a weighted tree, with the weight function taking values in an editable space, by calling it dendrogram, as formalized by the following definition.
@@
3.1 Finite Posets, DAGs and Transitive Reductions
In this section, we justify our focus on tree-shaped objects by outlining, at a high level, the various contexts in which such graphs arise in TDA. This also yields a conceptual blueprint for potential extensions of the present work.
We start from the definition of a directed graph.
@@ 3.1 Finite Posets, DAGs and Transitive Reductions
A directed graph is made by a pair $G=(V,E)$ with $V$ being the set of its vertices and $E\subseteq V\times V$ the set of its directed edges.
Consider a finite partially ordered set $(P,\leq)$. It is well known that one can associate to $P$ a unique (up to directed graph isomorphism) directed acyclic graph (DAG) $G_{P}$ so that $G_{P}$ is in bijection with the objects in $P$ and $(a,b)\in E_{G_{P}}$ if and only if $a\leq b$. Moreover, the graph $G_{P}$ can be simplified without losing ordering information. We know formalize these constructions.
Definition 7 (Aho et al. (1972))
A directed graph $G=(V,E)$ is called transitive if $(v,v^{\prime})\in E$, whenever there is a directed path from $v$ to $v^{\prime}$. For every directed graph $G=(V,E)$ we define its transitive closure as the graph $G^{T}:=(V,E_{G^{T}})$, with:
$E_{G^{T}}=\bigcap_{\{E^{\prime}\subset V\times V\mid E\subset E^{\prime},(V,E^{\prime})\text{ transitive graph}\}}E^{\prime}.$
@@
$E_{G^{t}}:=\bigcap_{\{E^{\prime}\subset V\times V\mid(V,E^{\prime})^{T}=G^{T}\}}E^{\prime}.$
Given a finite poset $P$, we call $\mathcal{G}(P)$ the transitive reduction of the associated DAG.
3.2 From Persistent Sets to Dendrograms
We now establish a connection between functors, posets, and tree. To this end, we introduce a general procedure that applies beyond the specific case of functors $S:(\mathbb{R},\leq)\rightarrow\operatorname*{Sets}$, called persistent sets (Carlsson and Mémoli, 2013; Curry, 2018; Curry et al., 2022), which constitutes the primary focus of this work. Such objects are in fact of great interest in TDA and represent a (non-equivalent) alternative to persistence modules/diagrams when working with information collected from path-connected components of filtrations. A more thorough exploration of the broader scenarios is left for future research.
The objects we introduce in the following are represented in Figure 2.